A negatively charged cart moves along a frictionless track from point A to point B. The cart has enough speed to make it all the way to point B. There is a non-constant electric field along the track. The field values change at each meter of the track and are known ( Toward B). This info is given below. At which meter marker is the cart moving the fastest? Explain

ITHAT THE E. He started in 1991. seen as a FREE 1 m 2 m 3 m 4 m 5 m 6 m 7 m 8 m 9 mn 10 m distance [m] 150 50 o .50 100 200 1100 a 50 0 10g E field Vm]
Answer #1

initial speed is v.

F = qE

and a = F/m = qE/m

so acceleration will be in opposite direction of E. (as q is negative)

upto 2 m , a will be in opposite direction of velocity .

after 4 to 8m , acceleration is in the direttion of velocity.

so a > 0

hence speed will increase upto 8m then it will be constant upto 9m

then speed will start to decrease.

hence speed is maximum at 8m to 9m.